# From Static to Adaptive Defense: Federated Multi-Agent Deep Reinforcement Learning-Driven Moving Target Defense Against DoS Attacks in UAV Swarm Networks

### v1

@@ Abstract
The proliferation of unmanned aerial vehicle (UAV) swarms has enabled a wide range of mission-critical applications, but also exposes UAV networks to severe Denial-of-Service (DoS) threats due to their open wireless environment, dynamic topology, and resource constraints. Traditional static or centralized defense mechanisms are often inadequate for such dynamic and distributed scenarios. To address these challenges, we propose a novel federated multi-agent deep reinforcement learning (FMADRL)-driven moving target defense (MTD) framework for proactive and adaptive DoS mitigation in UAV swarm networks. Specifically, we design three lightweight and coordinated MTD mechanisms, including leader switching, route mutation, and frequency hopping, that leverage the inherent flexibility of UAV swarms to disrupt attacker efforts and enhance network resilience. The defense problem is formulated as a multi-agent partially observable Markov decision process (POMDP), capturing the distributed, resource-constrained, and uncertain nature of UAV swarms under attack. Each UAV is equipped with a local policy agent that autonomously selects MTD actions based on partial observations and local experiences. By employing a policy gradient-based FMADRL algorithm, UAVs collaboratively optimize their defense policies via reward-weighted aggregation, enabling distributed learning without sharing raw data and thus reducing communication overhead. Extensive simulations demonstrate that our approach significantly outperforms state-of-the-art baselines, achieving up to a 34.6% improvement in attack mitigation rate, a reduction in average recovery time of up to 94.6%, and decreases in energy consumption and defense cost by as much as 29.3% and 98.3%, respectively, while maintaining robust mission continuity under various DoS attack strategies.
@@
Unmanned aerial vehicle swarm network, denial-of-service attacks, moving target defense, federated multi-agent deep reinforcement learning, policy gradient method.
@@ I Introduction
The rapid development of the unmanned aerial vehicle (UAV) technology [1] has enabled a wide range of applications, including environmental monitoring, disaster response, precision agriculture, logistics, aerial photography, and intelligent surveillance. By leveraging the collaborative capabilities of multiple UAVs, the UAV swarm [2] can achieve enhanced coverage, resilience, and real-time data processing, making them indispensable in both civilian and industrial domains. As the low-altitude economy continues to expand, UAV swarm networks [3] are expected to play an increasingly important role in smart cities, emergency management, and next-generation communication infrastructures.
@@
Nevertheless, the widespread adoption of UAV swarms also brings new security challenges [4]. Due to their reliance on open wireless links, limited energy and processing capabilities, UAV networks are particularly vulnerable to Denial-of-Service (DoS) attacks [5, 6]. For example, UAVs often operate without robust authentication or traffic filtering mechanisms. Attackers can easily launch attacks by overwhelming communication channels or computational resources, leading to service disruptions or UAV disconnection from the swarm [7]. In mission-critical scenarios, even a brief loss of connectivity or control can have catastrophic consequences, including the failure of time-sensitive missions or the crash of drones.
@@
In addition, UAV swarms operate in highly dynamic, resource-constrained, and often uncertain environments. The mobility of UAVs, the need for low-latency communication, and the distributed nature of control introduce unique challenges for both attack detection and defense[8]. Traditional security mechanisms, such as firewalls [9], intrusion detection systems (IDSs) [10], and traffic redirection, typically rely on prior knowledge of attack characteristics and are often designed for centralized infrastructures, making them ill-suited for UAV networks. Moreover, the static nature of these defenses results in several shortcomings: (i) limited adaptability to evolving and sophisticated attack patterns, (ii) increased false positive rates due to lack of contextual awareness, and (iii) reactive mitigation that only occurs after attacks have already caused damage. Therefore, ensuring the resilience of UAV networks against DoS attacks by more proactive and adaptive defense strategies is a critical research challenge.
@@
Fortunately, Moving Target Defense (MTD) [11, 12] has emerged as a promising approach for proactively defending against various cyber threats. MTD aims to increase the uncertainty for attackers and disrupt attack reconnaissance-effort asymmetry by continuously or periodically changing the attack surface of a system [13]. However, current MTD implementations face three critical limitations in UAV contexts. (i) Existing MTD mechanisms are not primarily proposed for UAV networks, and the overhead of frequent adaptations may exceed the limited resources available on UAV platforms. (ii) Most MTD decision-making frameworks are designed to collect global information and coordinate defense actions in a centralized manner, which cannot operate efficiently and reliably within the unique constraints of UAV swarm networks. (iii) Sophisticated attackers can reconstruct attack target states by tracking and learning the mutation patterns, enabling them to persistently target UAV networks despite the use of MTD. These limitations highlight the urgent need for distributed and intelligent MTD framework that can adaptively balance security and performance in UAV scenarios.
@@
In this paper, we propose a novel federated multi-agent deep reinforcement learning (FMADRL)-driven MTD framework tailored for UAV swarm networks facing DoS attacks, where the bird’s-eye view of the proposed method has been illustrated in Fig. 1. In our framework, we design three lightweight MTD actions, including (i) Leader Switching, (ii) Route Mutation, and (iii) Frequency Hopping, based on partial observations. The first mechanism allows the UAV swarm to promptly reassign the leader role among eligible UAVs when the current leader is persistently targeted, while the second mechanism dynamically reconfigures communication paths by selecting alternative relay UAVs, ensuring that critical messages can bypass compromised links and reach their destinations. Furthermore, frequency hopping periodically changes the communication frequency channels used by the swarm, significantly increasing the uncertainty for attackers and disrupting their ability to sustain effective attacks over time. To effectively mitigate adaptive DoS attacks while minimizing resource consumption and maintaining mission continuity, we first formulate the defense problem as a multi-agent partially observable Markov decision process (POMDP) and develop a policy gradient-based FMADRL (PG-FMADRL) method, where each UAV is treated as an independent agent that can learn and adapt its defense strategies based on local observations and experiences. Then, each agent periodically shares only its model parameters with a central aggregator using a federated learning scheme. This collaborative approach enables the UAV swarm to adaptively coordinate distributed defense strategies in real time without sharing raw data, thus respecting the stringent resource and latency constraints of UAV networks. The main contributions of this paper are summarized as follows:
@@ II-A DoS Attacks Detection and Mitigation in UAV Networks
DoS attacks in UAV networks have been extensively studied, with various detection and mitigation techniques proposed. For instance, Fu et al. [14] proposed an IDS for UAV networks, integrating convolutional neural network (CNN) and long short-term memory (LSTM) to achieve high detection accuracy of 94.4% for DoS attacks. Similarly, Hassler et al. [15] employed a method that fuses UAV cyber and physical features, which achieves an accuracy of at most 98.5% for attacks. To overcome challenges such as small sample sizes and uneven data distribution among UAVs, He et al. [16] leveraged a conditional generative adversarial network (CGAN)-based intrusion detection algorithm, achieving an accuracy of 99% in detecting DoS attacks. However, these methods typically require large amounts of labeled data for training, which may not be available in real-world scenarios, especially when timely response to novel or unseen attacks is required. Unlike existing methods, our approach operates without relying on prior knowledge or explicit attack signature identification, and instead dynamically orchestrates multiple MTD mechanisms to proactively respond to attacks in real time.
In addition to detection, several mitigation strategies have been proposed. For example, Gupta et al. [17] developed a machine learning (ML)-based Distributed DoS (DDoS) mitigation framework that leverages SDN’s programmability and centralized control to enable intelligent traffic management in UAV environments. Recent research has also investigated adaptive control strategies to enhance the resilience of UAV swarms against DoS attacks. Wu et al. [18] proposed a zero-sum differential game approach to effectively handle connectivity disruptions caused by DoS attacks and ensure boundedness and consensus control performance among UAVs. Tang et al. [19] incorporated robustness constraints to enhance disturbance resilience and introduced a dynamic event-triggered mechanism to respond to varying attack durations. Nevertheless, these approaches usually enhance the robustness of UAV swarms through tolerance and isolation strategies, which may be insufficient against sophisticated adversaries. To cope with this challenge, our proposed method fully leverages MTD techniques to present dynamic external attributes for UAV swarms, and enables them to autonomously adjust their defense strategies according to real-time network conditions and attack patterns.
@@ II-B MTD-based Solutions for DoS/DDoS Attacks
Recently, various MTD-based approaches have been introduced to counter DoS/DDoS attacks [20, 21, 22]. These methods aim to increase the attackers’ effort and cost by misleading them toward incorrect targets, thereby effectively diverting attacks away from the protected system. For example, Ribeiro et al. [23] proposed a novl architecture leveraging Software Defined Networking (SDN) for flow classification and MTD techniques for DDoS mitigation. Similarly, Zhang et al. [24] developed a collaborative mutation-based MTD (CM-MTD) framework to disrupt attacks. The authors formulated the MTD deployment as a semi-Markov decision process (SMDP), leveraging a hierarchical deep reinforcement learning (DRL) algorithm for scheduling of MTD actions. However, these MTD mechanisms are equipped with high computation resources, which are not suitable for resource-constrained environments such as UAV networks. In this study, we repurpose existing functionalities within UAVs as lightweight MTD mechanisms, enabling seamless integration into the dynamic defense of UAV swarms, while maintaining high effectiveness against DoS attacks.
To enable DoS/DDoS defense in resource-constrained environments, several recent studies have explored lightweight MTD mechanisms. For instance, Zhang et al. [25] designed configuration mutation mechanisms against DDoS attacks, where the communication ranges and capacities of roadside units (RSUs) are dynamically adjusted using a DRL algorithm. Similarly, Zhou et al. [26] introduced lightweight MTD mechanisms for the Internet of Things (IoT) environments, focusing on the dynamic control of device admission and service replica migration to defeat DDoS attacks. In the edge cloud context, the authors [13] designed several container-based MTD mechanisms and employed a deep Q-network (DQN) algorithm to optimize the trade-off between security and overhead when mitigating DDoS attacks. However, these mechanisms are not directly applicable to DoS defense in UAV networks. Furthermore, most existing studies that employ DRL for MTD optimization adopt centralized decision-making paradigms, which are ill-suited for the distributed and dynamic nature of UAV swarms. In contrast, our work proposes novel MTD mechanisms specifically designed for UAV environments and leverages a federated learning architecture to facilitate adaptive and collaborative defense across UAV swarm networks.
@@ III-A System Model
We consider a multi-UAV swarm system composed of a ground control station (GCS) and $N$ UAVs, denoted as $\mathcal{V}=\{0,1,...,N\}$, where node 00 is the GCS and nodes $1$ to $N$ are UAVs. The system operates in a three-dimensional space and is tasked with persistent patrol and robust formation maintenance.
@@
The communication network among the UAVs and the GCS is modeled as a dynamic undirected graph $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t))$. An edge $(i,j)\in\mathcal{E}(t)$ exists if and only if UAVs $i$ and $j$ are within a communication range $R_{c}$, operate on the same frequency channel $f$ at time $t$. Formally,
@@ III-B2 Attacker Strategies
  • Random Attacker: At each attack opportunity, the attacker randomly selects a UAV node or communication link from the set of available targets [30]. The attacker launches an attack on the chosen target for a fixed attack duration $\tau_{\text{atk}}$. After each attack, the attacker needs to wait for a reconnaissance period $\tau_{\text{recon}}$ before initiating the next attack. For example, if the random attacker selects a UAV $i$ as the target at time step 00, then the attack can be described as $\phi^{\mathcal{N}}_{i}(t)=1$ and $f_{\text{atk}}(t)=f_{i}(0)$ for $t\in[0,\tau_{\text{atk}}+\tau_{\text{recon}})$, and $\phi^{\mathcal{N}}_{j}(t)=0$ for all other UAVs $j\neq i$. Thus, the next attack can only start after initial attack ends and the attack frequency is updated to match the current frequency of the another selected UAV $j$, i.e., $f_{\text{atk}}(\tau_{\text{atk}}+\tau_{\text{recon}})=f_{j}(\tau_{\text{atk}}+\tau_{\text{recon}})$.
  • Greedy Attacker: The adversary dynamically observes the network state, enabling adaptively selecting the target that is expected to cause the maximum disruption to the swarm [31]. For example, the greedy attacker can select the current leader UAV $\mathbb{L}^{\mathcal{N}}(t)$ or the core link $\mathbb{L}^{\mathcal{L}}(t)$ of the swam as the target at time step $t$. For the sake of not weakening the greedy attacker’s ability, we assume that he/she may be aware of the existence of defense mechanisms (e.g., MTD), but not their specific deployment or timing. Therefore, the attacker can re-evaluate and update his/her target at time step $t^{\prime}$ to maximize attack effectiveness, but attack actions are subject to duration and reconnaissance constraints that are similar to those of the random attacker, i.e., $t^{\prime}=t+\tau_{\text{atk}}+\tau_{\text{recon}}$.
@@
where $\tau_{\text{eff}}=\tau_{\text{atk}}+\tau_{\text{recon}}$ is the lasting time of a round of attack. Similarly, the link attack can be described as

### v2

@@ Abstract
The proliferation of unmanned aerial vehicles (UAVs) has enabled a wide range of mission-critical applications and is becoming a cornerstone of low-altitude networks, supporting smart cities, emergency response, and more. However, the open wireless environment, dynamic topology, and resource constraints of UAVs expose low-altitude networks to severe Denial-of-Service (DoS) threats, undermining their reliability and security. Traditional defense approaches, which rely on fixed configurations or centralized decision-making, cannot effectively respond to the rapidly changing conditions in UAV swarm environments. To address these challenges, we propose a novel federated multi-agent deep reinforcement learning (FMADRL)-driven moving target defense (MTD) framework for proactive DoS mitigation in low-altitude networks. Specifically, we design lightweight and coordinated MTD mechanisms, including leader switching, route mutation, and frequency hopping, to disrupt attacker efforts and enhance network resilience. The defense problem is formulated as a multi-agent partially observable Markov decision process (POMDP), capturing the uncertain nature of UAV swarms under attack. Each UAV is equipped with a policy agent that autonomously selects MTD actions based on partial observations and local experiences. By employing a policy gradient-based FMADRL algorithm, UAVs collaboratively optimize their policies via reward-weighted aggregation, enabling distributed learning without sharing raw data and thus reducing communication overhead. Extensive simulations demonstrate that our approach significantly outperforms state-of-the-art baselines, achieving up to a 34.6% improvement in attack mitigation rate, a reduction in average recovery time of up to 94.6%, and decreases in energy consumption and defense cost by as much as 29.3% and 98.3%, respectively, under various DoS attack strategies. These results highlight the potential of intelligent, distributed defense mechanisms to protect low-altitude networks, paving the way for reliable and scalable low-altitude economy.
@@
Low-altitude networks, unmanned aerial vehicles, denial-of-service attacks, moving target defense, federated multi-agent deep reinforcement learning.
@@ I Introduction
The rapid development of unmanned aerial vehicle (UAV) technology [1] has enabled a wide range of applications, including environmental monitoring, disaster response, precision agriculture, logistics, aerial photography, and intelligent surveillance [2]. By leveraging the collaborative capabilities of multiple UAVs [3], the UAV swarm [4, 5] can achieve enhanced coverage, resilience, and real-time data processing [6], making them indispensable in both civilian and industrial domains. It is expected to play an increasingly important role in smart cities, emergency management, and next-generation communication infrastructures, forming the backbone of low-altitude networks.
@@
Nevertheless, the widespread adoption of UAV swarms also brings new security challenges [7] to low-altitude networks. Due to their reliance on open wireless links, limited energy and processing capabilities, UAV networks are particularly vulnerable to Denial-of-Service (DoS) attacks [8, 9]. For example, UAVs often operate without robust authentication or traffic filtering mechanisms. Attackers can easily launch attacks by overwhelming low-altitude communication channels or computational resources, leading to service disruptions or UAV disconnection from the swarm [10]. In mission-critical scenarios, even a brief loss of connectivity or control can have catastrophic consequences, including the failure of time-sensitive missions or the crash of drones.
@@
In addition, UAV swarms operate in highly dynamic, resource-constrained, and often uncertain environments. The mobility of UAVs, the need for low-latency communication, and the distributed nature of control introduce unique challenges for both attack detection and defense[11, 12]. Traditional security mechanisms, such as firewalls [13], intrusion detection systems (IDSs) [14], and traffic redirection, typically rely on prior knowledge of attack characteristics and are often designed for centralized infrastructures, making them ill-suited for low-altitude networks [15]. Moreover, the static nature of these defenses results in several shortcomings: (i) limited adaptability to evolving and sophisticated attack patterns, (ii) increased false positive rates due to lack of contextual awareness, and (iii) reactive mitigation that only occurs after attacks have already caused damage. Therefore, ensuring the resilience of low-altitude networks against DoS attacks by more proactive and adaptive defense strategies is a critical challenge.
To address this, Moving Target Defense (MTD) [16, 17] has emerged as a promising security paradigm that aims to increase the complexity and cost for attackers by continuously or periodically changing the system’s attack surface [18]. Instead of relying on static configurations, MTD introduces dynamic and unpredictable changes, such as IP randomization and port hopping, so that attackers cannot easily gather information or exploit vulnerabilities. This approach fundamentally shifts the traditional defender-attacker asymmetry, making it significantly harder for adversaries to plan and execute successful attacks.
@@
However, applying traditional MTD to UAV scenarios presents three critical challenges. (i) First, many MTD mechanisms consume significant resources, making them unsuitable for resource-constrained UAV platforms. (ii) Second, most MTD decision-making frameworks are designed to collect global information and coordinate defense actions in a centralized manner, which cannot operate efficiently and reliably within the unique constraints of low-altitude networks. (iii) Third, sophisticated attackers can reconstruct attack target states by tracking and learning the mutation patterns, enabling them to persistently target UAVs despite the use of MTD. These limitations highlight the urgent need for distributed and intelligent MTD framework that can adaptively balance security and performance in UAV scenarios.
@@
In this paper, we propose a novel federated multi-agent deep reinforcement learning (FMADRL)-driven MTD framework tailored for UAV swarm networks facing DoS attacks, where the bird’s-eye view of the proposed method has been illustrated in Fig. 1. In our framework, we design three lightweight MTD actions, including (i) Leader Switching, (ii) Route Mutation, and (iii) Frequency Hopping, based on partial observations. The first mechanism allows the UAV swarm to promptly reassign the leader role among eligible UAVs when the current leader is persistently targeted, while the second mechanism dynamically reconfigures communication paths by selecting alternative relay UAVs, ensuring that critical messages can bypass compromised links and reach their destinations. Furthermore, frequency hopping periodically changes the communication frequency channels used by the swarm, significantly increasing the uncertainty for attackers and disrupting their ability to sustain effective attacks over time. To effectively mitigate adaptive DoS attacks while minimizing resource consumption and maintaining mission continuity, we first formulate the defense problem as a multi-agent partially observable Markov decision process (POMDP) and develop a policy gradient-based FMADRL (PG-FMADRL) method, where each UAV is treated as an independent agent that can learn and adapt its defense strategies based on local observations and experiences. Then, each agent periodically shares only its model parameters with a central aggregator using a federated learning scheme. This collaborative approach enables the UAV swarm to adaptively coordinate distributed defense strategies in real time without sharing raw data, thus respecting the stringent resource and latency constraints of low-altitude networks. The main contributions of this paper are summarized as follows:
@@ II Related Work
This section reviews relevant literature on DoS attack mitigation in UAV networks and MTD-based defense mechanisms. We identify limitations in existing approaches and highlight the research gap our work addresses.
@@ II-A DoS Attacks Detection and Mitigation in UAV Networks
DoS attacks in UAV networks have been extensively studied, with various detection and mitigation techniques proposed. For instance, Fu et al. [19] proposed an IDS for UAV networks, integrating convolutional neural network (CNN) and long short-term memory (LSTM) to achieve high detection accuracy of 94.4% for DoS attacks. Similarly, Hassler et al. [20] employed a method that fuses UAV cyber and physical features, which achieves an accuracy of up to 98.5% for attacks. To overcome challenges such as small sample sizes and uneven data distribution among UAVs, He et al. [21] leveraged a conditional generative adversarial network (CGAN)-based intrusion detection algorithm, achieving an accuracy of 99% in detecting DoS attacks. However, these methods typically require large amounts of labeled data for training, which may not be available in real-world scenarios, especially when timely response to novel or unseen attacks is required.
In addition to detection, several mitigation strategies have been proposed. For example, Gupta et al. [22] developed a machine learning (ML)-based Distributed DoS (DDoS) mitigation framework that leverages SDN’s programmability and centralized control to enable intelligent traffic management in UAV environments. Recent research has also investigated adaptive control strategies to enhance the resilience of UAV swarms against DoS attacks. Wu et al. [23] proposed a zero-sum differential game approach to effectively handle connectivity disruptions caused by DoS attacks and ensure boundedness and consensus control performance among UAVs. Tang et al. [24] incorporated robustness constraints to enhance disturbance resilience and introduced a dynamic event-triggered mechanism to respond to varying attack durations. Nevertheless, these approaches usually enhance the robustness of UAV swarms through tolerance and isolation strategies, which may be insufficient against sophisticated adversaries.
Compared with the above methods, our approach offers several significant advantages. First, our method does not rely on prior knowledge, labeled data, or attack signatures, making it more robust to novel and adaptive threats. Second, by orchestrating multiple lightweight MTD mechanisms, our framework can rapidly respond to attacks in real time, minimizing service disruption and resource consumption. Third, the federated multi-agent reinforcement learning design enables collaborative defense without sharing raw data, which is particularly suitable for resource-constrained environments. These features collectively distinguish our work from prior studies and address key challenges in securing low-altitude networks.
@@ II-B MTD-based Solutions for DoS/DDoS Attacks
The core idea of MTD is to reduce the predictability and exposure of critical system components by introducing dynamic changes to the system’s configuration or behavior. Therefore, typical MTD techniques are designed to invalidate the attacker’s reconnaissance efforts, force them to adapt continuously, and ultimately increase the likelihood of attack failure. In recent years, MTD has been widely studied to counter DoS/DDoS attacks in both traditional and emerging network environments [25, 26, 27].
For example, Ribeiro et al. [28] proposed a novel architecture leveraging Software Defined Networking (SDN) for flow classification and MTD techniques for DDoS mitigation. Similarly, Zhang et al. [29] developed a collaborative mutation-based MTD (CM-MTD) framework to disrupt attacks. The authors formulated the MTD deployment as a semi-Markov decision process (SMDP), leveraging a hierarchical deep reinforcement learning (DRL) algorithm for scheduling of MTD actions. However, these MTD mechanisms are equipped with high computation resources, which are not suitable for resource-constrained environments such as UAV networks.
To enable DoS/DDoS defense in resource-constrained environments, several recent studies have explored lightweight MTD mechanisms. For instance, Zhang et al. [30] designed configuration mutation mechanisms against DDoS attacks, where the communication ranges and capacities of roadside units (RSUs) are dynamically adjusted using a DRL algorithm. Similarly, Zhou et al. [31] introduced lightweight MTD mechanisms for the Internet of Things (IoT) environments, focusing on the dynamic control of device admission and service replica migration to defeat DDoS attacks. In the edge cloud context, the authors [18] designed several container-based MTD mechanisms and employed a deep Q-network (DQN) algorithm to optimize the trade-off between security and overhead when mitigating DDoS attacks. However, these mechanisms are not directly applicable to DoS defense in UAV networks. Furthermore, most existing studies that employ DRL for MTD optimization adopt centralized decision-making paradigms, which are ill-suited for the distributed and dynamic nature of UAV swarms.
In summary, while prior MTD-based solutions have demonstrated effectiveness in specific scenarios, our work is the first to propose a federated and multi-agent MTD framework specifically tailored for low-altitude networks. Unlike previous studies that rely on centralized decision-making or heavyweight mutations, our approach enables distributed, adaptive, and energy-efficient defense by leveraging lightweight mechanisms and federated learning. Furthermore, our method optimizes MTD deployment through intelligent policy learning, achieving superior robustness and scalability in dynamic, low-altitude network scenarios. These unique advantages highlight the novelty and practical value of our contribution.
@@ III System Model and Problem Formulation
This section presents the system architecture and threat model that form the foundation of our work. We first introduce the UAV swarm model, then define the DoS attack types and attacker strategies, followed by our proposed MTD mechanisms. Finally, we formulate the defense problem as a multi-objective optimization problem and analyze its characteristics.
@@ III-A System Model
We consider a multi-UAV swarm system composed of a ground control station (GCS) and $N$ UAVs, denoted as $\mathcal{V}=\{0,1,...,N\}$, where node 0 is the GCS and nodes $1$ to $N$ are UAVs. The system operates in a three-dimensional space and is tasked with persistent patrol and robust formation maintenance.
@@
To understand communication possibilities within the swarm, we represent the network as a dynamic undirected graph $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t))$. An edge $(i,j)\in\mathcal{E}(t)$ exists if and only if UAVs $i$ and $j$ are within a communication range $R_{c}$, operate on the same frequency channel $f$ at time $t$. Formally,
@@ III-B2 Attacker Strategies
  • Random Attacker: At each attack opportunity, the attacker randomly selects a UAV node or communication link from the set of available targets [35]. The attacker launches an attack on the chosen target for a fixed attack duration $\tau_{\text{atk}}$. After each attack, the attacker needs to wait for a reconnaissance period $\tau_{\text{recon}}$ before initiating the next attack. During this reconnaissance period, the attacker is assumed to be passively listening or actively scanning to re-acquire a target, by identifying the swarm’s current communication frequency if a frequency hop has occurred. For example, if the random attacker selects a UAV $i$ as the target at time step 0, then the attack can be described as $\phi^{\mathcal{N}}_{i}(t)=1$ and $f_{\text{atk}}(t)=f_{i}(0)$ for $t\in[0,\tau_{\text{atk}}+\tau_{\text{recon}})$, and $\phi^{\mathcal{N}}_{j}(t)=0$ for all other UAVs $j\neq i$. Thus, the next attack can only start after initial attack ends and the attack frequency is updated to match the current frequency of the another selected UAV $j$, i.e., $f_{\text{atk}}(\tau_{\text{atk}}+\tau_{\text{recon}})=f_{j}(\tau_{\text{atk}}+\tau_{\text{recon}})$.
  • Greedy Attacker: The adversary dynamically observes the network state, enabling adaptively selecting the target that is expected to cause the maximum disruption to the swarm [36]. For example, the greedy attacker can select the current leader UAV $\mathbb{L}^{\mathcal{N}}(t)$ or the core link $\mathbb{L}^{\mathcal{L}}(t)$ of the swam as the target at time step $t$. For the sake of not weakening the greedy attacker’s ability, we assume that he/she may be aware of the existence of defense mechanisms (e.g., MTD), but not their specific deployment or timing. Therefore, the attacker can re-evaluate and update his/her target at time step $t^{\prime}$ to maximize attack effectiveness, but attack actions are subject to duration and reconnaissance constraints that are similar to those of the random attacker, i.e., $t^{\prime}=t+\tau_{\text{atk}}+\tau_{\text{recon}}$. This reconnaissance period represents the time required for the adversary to gather new intelligence, such as identifying the new leader after a switch or discovering the new communication channel after a frequency hop, before launching the next targeted attack.
@@
where we define the effective attack duration as $\tau_{\text{eff}}=\tau_{\text{atk}}+\tau_{\text{recon}}$, representing the total time for one round of attack and reconnaissance. Similarly, the link attack can be described as